Category Archives: Observing

Here we update what the current status of the observing period is.

The week of February 27

This week marked the end of another dark run and thus the end of another science period. We are now also three quarters of the way through the current science period labeled 17-1. As such we are starting to see a larger number of filler Priority 4 targets being observed. In this last dark run we observed 374 acceptable observations for a total of 38 hours and 122 Priority 4 targets for a total of 22 hours. We saw 6% of the time lost due to problems. This is a fairly respectable number and fairly typical of the mature HET before the takedown.

We are now three quarters of the way through the current 17-1 period. The Telescope Allocation Committees have begun meeting and the new allocations are being submitted for the 17-2 period which begins April 1. For this period the same suite of instruments will be available for the PIs: LRS2-B and LRS2-R and the two weeks around dark time will also be fully committed to science observing.

The Week of January 30th

The end of this week marks us as half way through the current 4 month period (December – March). We have completed 70% of the highest priority band P1, 52% of the P2 band and 41% of the P3 band time. The bad weather we had at the beginning of the period, back in December, really kept us from completing many of the high priority targets in the Sheila field which is now unreachable until the late summer.

There are 10 programs that have been completed out of 27 programs. Some of our users have never been granted HET time before so there is a bit of a learning curve in how to optimize the HET time. Hopefully our Synoptic Supernova teams can make excellent use of the remaining 2 months of the period.

For the next 2 weeks (during the bright time) we will be concentrating on making further improvements to our telescope control systems, wavefront sensor systems and instrument control systems. It is going to be a very busy 2 weeks!

Update from January 9 2017

We have finished another dark period from the first trimester back observing with the HET. Recall that we are only observing during the two weeks around the new moon. Thus, we finished up last week with our targets that can be observed with some moon. We have completed 63.4 hours of charged time with typical setup times of 10 minutes for LRS2 targets. Some programs, particularly those from our smaller partners, have been completed finished while a few of the larger programs have only had a small dent made in their total Telescope Allocation Committee (TAC) allocations.

There are a few fields that seem to be very popular. The Shelia and Cosmos fields are among the most heavily over-subscribed. We will not be able to complete the requested visits to the Shelia fields but if the weather permits and the engineering efforts during the 1st quarter are not to heavy we might try to sneak in a few more targets.

We are developing a hole at the start of the night after the Shelia field sets and will likely have to deal with that in the next few weeks unless some new targets arrive.

Week of December 26th

It has been quiet during the day at the HET this week with day staff on holiday. If the weather clears the night staff will be plenty busy working through the queue. The current weather outlook for the week is cloudy with possible showers. Monday night we were able to do science through thin clouds and successfully collected data for many programs that do not require perfect seeing or zero cloud cover. The bulk of the science data collected in the beginning of the week has been with LRS2-B and LRS2-R. Fingers crossed the weather clears as we approach the new moon on Thursday.

Week of December 19th

This week has been an exciting week ramping up to science.  The Board of Visitor Staff Excellence Award Winners were announced by Director Armandroff and are as follows: Henry Cantu, Angela Davis, Steve Odewahn, and Trent Peterson.  Congratulations to the winners.

We are quickly moving in to science mode with the upcoming new moon.  Currently this week the first half of each night is being used primarily for science with the second half used for engineering due to the moon.  The weather looks to be clearing over the holiday weekend, so hopefully we will be able to collect some good data for the HET community!

The week of December 11

This week marks the first full week of the new trimester (December – March). During this period we will be observing for 2 weeks around the new moon with LRS2-B and LRS2-R and to a very limited extent the first 16 units of VIRUS for science projects from our partners. We have 27 separate projects from our 4 partners. During this first week of the new trimester we took data for 10 programs.

During the 2 weeks around the full moon we will be concentrating on engineering but we may if it does not impose any constraint on the engineering obtain a few science targets during grey and bright time.

It is good to be back online taking science data.

Week of September 19th

Big Thunderstorm

One of the many powerful thunderstorms that passed by us this month. This one produced hail just north of the HET.

This week we finally had a break in the stormy weather and started wave front sensor commissioning.  A team from Austin came out and installed an imager and wave front sensor in the IHMP (Input Head Mounting Plate) to calibrate all other wave front sensors on the telescope.  These devices once calibrated will help keep the tracker in perfect alignment with the mirror to optimize the observations we are taking for science.  This will be the main focus for the next month.  Hopefully the weather is more cooperative than it has been.  If you would like to see a time-lapse of some of the recent storms follow this link: https://www.youtube.com/watch?v=shRD2WmPWd4

The week of July 11

Last week we reported on a incident with the HET tracker that put us in a hard skew condition and shut us down. We had recovered from that skew condition but the software issues that led to the skew had continued. This week I am happy to report that we have not only installed additional skew sensors but we have found the source of the software issues. The software teams in Austin and West Texas were able to perform a series of experiments which allowed them to isolate the problem as coming from the ethernet card inside the dSpace box. This is a very specialized ethernet card but thankfully we were able to swap out a different card for it and found that all of our TCS communications problems, hexapod faults, and loading trajectory errors have gone away. We will look into getting a spare card and send the card in to see what kind of fault occurred in the card.

With the improved tracker performance we were able to start our early science operations. We took our first science targets on the night of the 16th. Unfortunately, we are now headed towards full moon which is not the ideal time to do science targets with LRS2. However, it does give us a chance to work out any inefficiencies in our PI interface, night operations interfaces and queue tools so that we can be fully ready once we get some dark time a few days after full moon.

The week of July 4th

Two weeks ago we had a fairly serious tracker issue. While recovering from a fairly typical tracker failure, the telescope operator requested that the breaks be set but instead the tracker moved in X and generated a hard skew. This hard skew actually caused the bearing to come out of the bearing block. This occurred the day that the blogger and most of the chief engineers were going to their summer conferences. Upon the return the engineers were able to determine that no permanent damage was done to the bearings or the tracker in general. The engineering team was able to move the tracker back into the block using the tracker motors. Unfortunately, the software issues that led to the skew are still present and have not yet been sorted out. This has severely limited our ability to run trajectories for engineering or the start of early science operations.

In the mean time work continues on VIRUS enclosure 2 and with re-coating of the primary mirror segments.

The week of May 23

In the previous week the HET Board of Directors met and were generally happy with the progress being made on the HET. They are very eager for science operations to begin and have authorized the start of Early Science Operations starting July 1st. This will mean that HET member institutions will be able to submit requests to their local Telescope Allocation Committee in the coming weeks. The amount of time available will grow with time but is likely to start with 1 week per dark period. The instrument that will be available for Early Science Operations will be the new Low Resolution Spectrograph (LRS2).

In this week the HET continued with VIRUS commissioning during bright time with the addition of 7 new VIRUS units. This brings the number of spectrographs units (and IFUs) to 16. We will use these during commissioning in the coming dark run but we can already see that several of them will need some further engineering while others are very good.

The mirror team continues to swap mirrors into the array this week they swapped out 4 mirrors. The array now has an average coating age of 454 days with a swap rate of 1.7 segments/week.